DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, May 7, 2003 Padelford C-36 3:45 pm

A classification of conformally invariant differential operators in odd dimensions

Spyros Alexakis (Princeton University)

A new family of conformally invariant differential operators on densities is constructed and a classification result is proven in odd dimensions. This result complements the work of Graham, Mason, Jenne and Sparling where they construct conformally invariant curved analogues of the powers of the Laplacian. Furthermore, we find how these new operators govern the transformation under conformal rescaling of certain scalar quantities that depend upon the Riemannian curvature tensor.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.